# Self-Enhancing Video Data Management System for Compositional Events with Large Language Models

### v1

@@ 7. Related work
Video analytics. Numerous video analytics systems have been developed to support a wide range of data management tasks (Daum et al., 2023, 2022; Romero et al., 2023; Kittivorawong et al., 2023; Kossmann et al., 2023). Query execution over videos typically involves running expensive ML models. Thus, many techniques have been proposed to accelerate query processing, including indexing (Kang et al., 2022; He et al., 2020; Hu et al., 2022), sampling (Moll et al., 2020; Bastani et al., 2020a; Bang et al., 2023), pre-filtering frames (Lu et al., 2018; He et al., 2021, 2023; Xarchakos and Koudas, 2019), reusing results (Xu et al., 2022), and building specialized models (Kang et al., 2019; Anderson et al., 2019). VOCAL-UDF can incorporate existing methods to optimize query execution.
@@
Compositional video query processing. VOCAL-UDF is most related to systems designed for compositional video queries (Bastani et al., 2020b; Chen et al., 2022a; Fu et al., 2019; Liu et al., 2019; Chunduri et al., 2022; Yadav and Curry, 2019; Mell et al., 2023; Zhang et al., 2023; Yu et al., 2024; Xu et al., 2022). However, these systems often require users to have a certain level of database expertise to manually construct compositional queries (Chen et al., 2022a; Fu et al., 2019; Liu et al., 2019; Xu et al., 2022; Yu et al., 2024) or to provide a few examples for the system to learn a query from (Zhang et al., 2023; Mell et al., 2023). In contrast, VOCAL-UDF leverages the advances in LLMs and enables users to express queries in natural language.
@@
LLMs with tools. LLMs are widely used to tackle challenging text tasks across a variety of applications (Chen et al., 2023a; Guo et al., 2023; Sharma et al., 2021; Pang et al., 2024; Kayali et al., 2024; Chen et al., 2021). By integrating external tools, LLMs can address even more complex reasoning tasks (Lu et al., 2023; Schick et al., 2023; Qin et al., 2024; Zhou et al., 2023; Ma et al., 2024), including vision tasks (Mitra et al., 2024; Choudhury et al., 2023; Gupta and Kembhavi, 2023; Surís et al., 2023; Shen et al., 2023). However, these systems generally rely on the availability of existing tools or modules. For instance, VisProg (Gupta and Kembhavi, 2023) and ProViQ (Choudhury et al., 2023) utilize LLMs to transform complex tasks into executable programs that invoke predefined tools. VOCAL-UDF also employs LLMs to handle compositional queries over videos but extends this capability by generating new UDFs. Several systems explore the capacity of LLMs to create new tools. LATM (Cai et al., 2023) generates reusable code snippets for natural language tasks, while GENOME (Chen et al., 2023c) generates and reuses code-based modules to solve visual tasks. VOCAL-UDF distinguishes itself as an end-to-end video data management system by supporting the generation of both program-based and distilled-model UDFs, which we show can significantly enhance query performance.
@@ Acknowledgements.
This work was funded in part by NSF award 2211133.References
@@
- SCH (2013) 2013. Debris in the deep: Using a 22-year video annotation database to survey marine litter in Monterey Canyon, central California, USA. Deep Sea Research Part I: Oceanographic Research Papers 79 (2013), 96–105.
@@
- Austin et al. (2021) Jacob Austin, Augustus Odena, Maxwell I. Nye, Maarten Bosma, Henryk Michalewski, David Dohan, Ellen Jiang, Carrie J. Cai, Michael Terry, Quoc V. Le, and Charles Sutton. 2021. Program Synthesis with Large Language Models. CoRR abs/2108.07732 (2021).
- Bai et al. (2023) Jinze Bai, Shuai Bai, Shusheng Yang, Shijie Wang, Sinan Tan, Peng Wang, Junyang Lin, Chang Zhou, and Jingren Zhou. 2023. Qwen-VL: A Versatile Vision-Language Model for Understanding, Localization, Text Reading, and Beyond. arXiv preprint arXiv:2308.12966 (2023).
@@
- Bubeck et al. (2023) Sébastien Bubeck, Varun Chandrasekaran, Ronen Eldan, Johannes Gehrke, Eric Horvitz, Ece Kamar, Peter Lee, Yin Tat Lee, Yuanzhi Li, Scott M. Lundberg, Harsha Nori, Hamid Palangi, Marco Túlio Ribeiro, and Yi Zhang. 2023. Sparks of Artificial General Intelligence: Early experiments with GPT-4. CoRR abs/2303.12712 (2023).
@@
- Cai et al. (2023) Tianle Cai, Xuezhi Wang, Tengyu Ma, Xinyun Chen, and Denny Zhou. 2023. Large Language Models as Tool Makers. CoRR abs/2305.17126 (2023).
- Chen et al. (2024) Boyuan Chen, Zhuo Xu, Sean Kirmani, Brian Ichter, Danny Driess, Pete Florence, Dorsa Sadigh, Leonidas J. Guibas, and Fei Xia. 2024. SpatialVLM: Endowing Vision-Language Models with Spatial Reasoning Capabilities. CoRR abs/2401.12168 (2024).
- Chen et al. (2022b) Bei Chen, Fengji Zhang, Anh Nguyen, Daoguang Zan, Zeqi Lin, Jian-Guang Lou, and Weizhu Chen. 2022b. CodeT: Code Generation with Generated Tests. arXiv:2207.10397
@@
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Pondé de Oliveira Pinto, Jared Kaplan, Harrison Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, Alex Ray, Raul Puri, Gretchen Krueger, Michael Petrov, Heidy Khlaaf, Girish Sastry, Pamela Mishkin, Brooke Chan, Scott Gray, Nick Ryder, Mikhail Pavlov, Alethea Power, Lukasz Kaiser, Mohammad Bavarian, Clemens Winter, Philippe Tillet, Felipe Petroski Such, Dave Cummings, Matthias Plappert, Fotios Chantzis, Elizabeth Barnes, Ariel Herbert-Voss, William Hebgen Guss, Alex Nichol, Alex Paino, Nikolas Tezak, Jie Tang, Igor Babuschkin, Suchir Balaji, Shantanu Jain, William Saunders, Christopher Hesse, Andrew N. Carr, Jan Leike, Joshua Achiam, Vedant Misra, Evan Morikawa, Alec Radford, Matthew Knight, Miles Brundage, Mira Murati, Katie Mayer, Peter Welinder, Bob McGrew, Dario Amodei, Sam McCandlish, Ilya Sutskever, and Wojciech Zaremba. 2021. Evaluating Large Language Models Trained on Code. CoRR abs/2107.03374 (2021).
- Chen et al. (2019) Vincent S. Chen, Paroma Varma, Ranjay Krishna, Michael S. Bernstein, Christopher Ré, and Li Fei-Fei. 2019. Scene Graph Prediction With Limited Labels. ICCV (2019).
- Chen et al. (2023b) Xinyun Chen, Maxwell Lin, Nathanael Schärli, and Denny Zhou. 2023b. Teaching Large Language Models to Self-Debug. arXiv:2304.05128
- Chen et al. (2022a) Yueting Chen, Nick Koudas, Xiaohui Yu, and Ziqiang Yu. 2022a. Spatial and Temporal Constrained Ranked Retrieval over Videos. PVLDB 15, 11 (2022), 3226–3239.
- Chen et al. (2023a) Zui Chen, Lei Cao, Sam Madden, Tim Kraska, Zeyuan Shang, Ju Fan, Nan Tang, Zihui Gu, Chunwei Liu, and Michael Cafarella. 2023a. SEED: Domain-Specific Data Curation With Large Language Models. arXiv:2310.00749
- Chen et al. (2023c) Zhenfang Chen, Rui Sun, Wenjun Liu, Yining Hong, and Chuang Gan. 2023c. GENOME: GenerativE Neuro-symbOlic visual reasoning by growing and reusing ModulEs. ArXiv abs/2311.04901 (2023).
@@
- Dellinger et al. (2021) Justin Dellinger, Carolyn Shores, Apryle Craig, Shannon Kachel, Michael Heithaus, William Ripple, and Aaron Wirsing. 2021. Predators reduce niche overlap between sympatric prey. Oikos (12 2021). https://doi.org/10.1111/oik.08628
@@
- Feng et al. (2021) Qi Feng, Vitaly Ablavsky, and Stan Sclaroff. 2021. CityFlow-NL: Tracking and Retrieval of Vehicles at City Scale by Natural Language Descriptions. CoRR abs/2101.04741 (2021).
@@
- Fu et al. (2019) Daniel Y. Fu, Will Crichton, James Hong, Xinwei Yao, Haotian Zhang, Anh Truong, Avanika Narayan, Maneesh Agrawala, Christopher Ré, and Kayvon Fatahalian. 2019. Rekall: Specifying Video Events using Compositions of Spatiotemporal Labels. arXiv preprint arXiv:1910.02993 (2019).
@@
- Hammer et al. (2020) Patrick Hammer, Tony Lofthouse, Enzo Fenoglio, Hugo Latapie, and Pei Wang. 2020. A Reasoning Based Model for Anomaly Detection in the Smart City Domain. In IntelliSys (AISC), Vol. 1251. 144–159.
@@
- He et al. (2023) Dong He, Jieyu Zhang, Maureen Daum, Alexander Ratner, and Magdalena Balazinska. 2023. MaskSearch: Querying Image Masks at Scale. arXiv preprint arXiv:2305.02375 (2023).
@@
- Hinton et al. (2015) Geoffrey E. Hinton, Oriol Vinyals, and Jeffrey Dean. 2015. Distilling the Knowledge in a Neural Network. CoRR abs/1503.02531 (2015).
@@
- Hsieh et al. (2023a) Cheng-Yu Hsieh, Sibei Chen, Chun-Liang Li, Yasuhisa Fujii, Alexander J. Ratner, Chen-Yu Lee, Ranjay Krishna, and Tomas Pfister. 2023a. Tool Documentation Enables Zero-Shot Tool-Usage with Large Language Models. ArXiv abs/2308.00675 (2023). https://api.semanticscholar.org/CorpusID:260351459
@@
- Kakkar et al. (2024) Gaurav Tarlok Kakkar, Jiashen Cao, Aubhro Sengupta, Joy Arulraj, and Hyesoon Kim. 2024. Hydro: Adaptive Query Processing of ML Queries. CoRR abs/2403.14902 (2024).
@@
- Kittivorawong et al. (2023) Chanwut Kittivorawong, Yongming Ge, Yousef Helal, and Alvin Cheung. 2023. Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations. CoRR abs/2308.03276 (2023).
@@
- Lahiri et al. (2023) Shuvendu K. Lahiri, Sarah Fakhoury, Aaditya Naik, Georgios Sakkas, Saikat Chakraborty, Madanlal Musuvathi, Piali Choudhury, Curtis von Veh, Jeevana Priya Inala, Chenglong Wang, and Jianfeng Gao. 2023. Interactive Code Generation via Test-Driven User-Intent Formalization. arXiv:2208.05950
@@
- Ma et al. (2024) Zixian Ma, Weikai Huang, Jieyu Zhang, Tanmay Gupta, and Ranjay Krishna. 2024. m&m’s: A Benchmark to Evaluate Tool-Use for multi-step multi-modal Tasks. arXiv preprint arXiv:2403.11085 (2024).
@@
- Mitra et al. (2024) Chancharik Mitra, Brandon Huang, Trevor Darrell, and Roei Herzig. 2024. Compositional Chain of Thought Prompting for Large Multimodal Models. In CVPR.
@@
- Qiao et al. (2024) Wei Qiao, Tushar Dogra, Otilia Stretcu, Yu-Han Lyu, Tiantian Fang, Dongjin Kwon, Chun-Ta Lu, Enming Luo, Yuan Wang, Chih-Chun Chia, et al. 2024. Scaling Up LLM Reviews for Google Ads Content Moderation. In Proceedings of the 17th ACM International Conference on Web Search and Data Mining. 1174–1175.
@@
- Romero et al. (2023) Francisco Romero, Caleb Winston, Johann Hauswald, Matei Zaharia, and Christos Kozyrakis. 2023. Zelda: Video Analytics using Vision-Language Models. CoRR abs/2305.03785 (2023).
@@
- Rozière et al. (2023) Baptiste Rozière, Jonas Gehring, Fabian Gloeckle, Sten Sootla, Itai Gat, Xiaoqing Ellen Tan, Yossi Adi, Jingyu Liu, Tal Remez, Jérémy Rapin, Artyom Kozhevnikov, Ivan Evtimov, Joanna Bitton, Manish Bhatt, Cristian Canton-Ferrer, Aaron Grattafiori, Wenhan Xiong, Alexandre Défossez, Jade Copet, Faisal Azhar, Hugo Touvron, Louis Martin, Nicolas Usunier, Thomas Scialom, and Gabriel Synnaeve. 2023. Code Llama: Open Foundation Models for Code. CoRR abs/2308.12950 (2023).
@@
- Toubal et al. (2024) Imad Eddine Toubal, Aditya Avinash, Neil Gordon Alldrin, Jan Dlabal, Wenlei Zhou, Enming Luo, Otilia Stretcu, Hao Xiong, Chun-Ta Lu, Howard Zhou, Ranjay Krishna, Ariel Fuxman, and Tom Duerig. 2024. Modeling Collaborator: Enabling Subjective Vision Classification With Minimal Human Effort via LLM Tool-Use. arXiv:2403.02626 [cs.CV]
- Wan et al. (2023) Zhongwei Wan, Xin Wang, et al. 2023. Efficient Large Language Models: A Survey. arXiv preprint arXiv:2312.03863 (2023).
@@
- Wang et al. (2024) Jiayu Wang, Yifei Ming, Zhenmei Shi, Vibhav Vineet, Xin Wang, and Neel Joshi. 2024. Is A Picture Worth A Thousand Words? Delving Into Spatial Reasoning for Vision Language Models. CoRR abs/2406.14852 (2024).
@@
- Wu et al. (2023) Qingyun Wu, Gagan Bansal, Jieyu Zhang, Yiran Wu, Beibin Li, Erkang Zhu, Li Jiang, Xiaoyun Zhang, Shaokun Zhang, Jiale Liu, Ahmed Hassan Awadallah, Ryen W White, Doug Burger, and Chi Wang. 2023. AutoGen: Enabling Next-Gen LLM Applications via Multi-Agent Conversation Framework. arXiv:2308.08155

### v2

@@ 7. Related work
Video analytics. Numerous video analytics systems have been developed to support a wide range of data management tasks (Daum et al., 2023, 2022; Romero et al., 2023; Kittivorawong et al., 2024; Kossmann et al., 2023). Query execution over videos typically involves running expensive ML models. Thus, many techniques have been proposed to accelerate query processing, including indexing (Kang et al., 2022; He et al., 2020; Hu et al., 2022), sampling (Moll et al., 2020; Bastani et al., 2020a; Bang et al., 2023), pre-filtering frames (Lu et al., 2018; He et al., 2021, 2024; Xarchakos and Koudas, 2019), reusing results (Xu et al., 2022), and building specialized models (Kang et al., 2019; Anderson et al., 2019). VOCAL-UDF can incorporate existing methods to optimize query execution.
@@
Compositional video query processing. VOCAL-UDF is most related to systems designed for compositional video queries (Bastani et al., 2020b; Chen et al., 2022; Fu et al., 2019; Liu et al., 2019; Chunduri et al., 2022; Yadav and Curry, 2019; Mell et al., 2023; Zhang et al., 2023b; Yu et al., 2024; Xu et al., 2022). However, these systems often require users to have a certain level of database expertise to manually construct compositional queries (Chen et al., 2022; Fu et al., 2019; Liu et al., 2019; Xu et al., 2022; Yu et al., 2024) or to provide examples to learn a query from (Zhang et al., 2023b; Mell et al., 2023). In contrast, VOCAL-UDF leverages advances in LLMs and allows expressing queries in NL.
@@
LLMs with tools. LLMs are widely used to tackle challenging text tasks across a variety of applications (Chen et al., 2024a; Guo et al., 2023; Sharma et al., 2021; Pang et al., 2024; Kayali et al., 2024; Chen et al., 2021; Wang et al., 2025a, b). By integrating external tools, LLMs can address even more complex reasoning tasks (Lu et al., 2023; Schick et al., 2023; Qin et al., 2024; Zhou et al., 2024; Ma et al., 2024), including vision tasks (Mitra et al., 2024; Choudhury et al., 2023; Gupta and Kembhavi, 2023; Surís et al., 2023; Shen et al., 2023). However, these systems generally rely on the availability of existing tools or modules. For instance, VisProg (Gupta and Kembhavi, 2023) and ProViQ (Choudhury et al., 2023) utilize LLMs to transform complex tasks into executable programs that invoke predefined tools. VOCAL-UDF also employs LLMs to handle compositional queries over videos but extends this capability by generating new UDFs. Several systems explore the capacity of LLMs to create new tools. LATM (Cai et al., 2024) generates reusable code snippets for NL tasks, while GENOME (Chen et al., 2024c) generates and reuses code-based modules to solve visual tasks. VOCAL-UDF distinguishes itself as an end-to-end VDBMS that generates both program-based and distilled-model UDFs which significantly enhances performance.
8. Limitations
VOCAL-UDF does not currently support relationship and attribute UDFs involving state changes over time, such as identifying a car “moving close” to a truck, which requires distance comparisons at different timestamps. As an approximation, multiple frame-level UDFs can represent each state separately; for instance, detecting a “far” relationship followed by a “near” one.
Materializing and reusing UDF results can be problematic when the definition of an attribute or relationship is not objective (e.g., “near”) and varies from query to query. To address this, VOCAL-UDF could allow users to access and manage generated UDF descriptions, implementations, and results, enabling them to dynamically enable or disable specific UDFs before issuing queries.
VOCAL-UDF’s performance is affected by its ability to semantically understand the user queries and available UDFs. We assume that predefined UDFs are relevant to the target domain and that an LLM is capable of disambiguating each UDF effectively. The performance of VOCAL-UDF thus depends significantly on the underlying LLM, which is why we used the latest LLMs (e.g., GPT-4o) to ensure high-quality semantic understanding. Furthermore, VOCAL-UDF currently supports queries with detailed and explicit descriptions. Extending support to vague queries is left for future work.
VOCAL-UDF leverages LLMs in various system components, making the quality of query results dependent on the underlying LLM performance. To enhance the reliability of LLM-generated answers, VOCAL-UDF employs syntax verification, generates multiple candidate UDFs, and leverages user annotations via active learning to select the best implementation. Introducing dummy UDFs further ensures that newly generated UDFs do not negatively impact query results. While users already provide annotations to guide UDF selection, VOCAL-UDF could be extended to incorporate additional human interventions, allowing users to access intermediate results for manual examination (Yang et al., 2023; Zhang et al., 2023a; Cho et al., 2024). Our experiments empirically demonstrate the promising potential of our LLM-based approach, and we posit that as LLMs advance, VOCAL-UDF performance will correspondingly improve.
@@ Acknowledgements.
This work was funded in part by NSF award 2211133. Nicole Sullivan was supported in part by an NSF graduate fellowship (DGE-2140004).References
@@
- aws (2024a) 2024a. Amazon Bedrock Pricing. https://aws.amazon.com/bedrock/pricing.
- aws (2024b) 2024b. Amazon EC2 Spot Instances Pricing. https://aws.amazon.com/ec2/spot/pricing.
@@
- lla (2024) 2024. Llama 3 70B vs GPT-4: Comparison Analysis. https://www.vellum.ai/blog/llama-3-70b-vs-gpt-4-comparison-analysis.
- nvi (2024) 2024. NVIDIA DALI. https://developer.nvidia.com/dali.
- ope (2024) 2024. OpenAI API pricing. https://openai.com/api/pricing.
@@
- Austin et al. (2021) Jacob Austin, Augustus Odena, Maxwell I. Nye, Maarten Bosma, Henryk Michalewski, David Dohan, Ellen Jiang, Carrie J. Cai, Michael Terry, Quoc V. Le, and Charles Sutton. 2021. Program Synthesis with Large Language Models. arXiv:2108.07732 [cs.PL]
- Bai et al. (2023) Jinze Bai, Shuai Bai, Shusheng Yang, Shijie Wang, Sinan Tan, Peng Wang, Junyang Lin, Chang Zhou, and Jingren Zhou. 2023. Qwen-VL: A Versatile Vision-Language Model for Understanding, Localization, Text Reading, and Beyond. arXiv:2308.12966 [cs.CV]
@@
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. In NeurIPS.
- Bubeck et al. (2023) Sébastien Bubeck, Varun Chandrasekaran, Ronen Eldan, Johannes Gehrke, Eric Horvitz, Ece Kamar, Peter Lee, Yin Tat Lee, Yuanzhi Li, Scott M. Lundberg, Harsha Nori, Hamid Palangi, Marco Túlio Ribeiro, and Yi Zhang. 2023. Sparks of Artificial General Intelligence: Early experiments with GPT-4. arXiv:2303.12712 [cs.CL]
@@
- Cai et al. (2024) Tianle Cai, Xuezhi Wang, Tengyu Ma, Xinyun Chen, and Denny Zhou. 2024. Large Language Models as Tool Makers. In ICLR.
- Chen et al. (2024d) Boyuan Chen, Zhuo Xu, Sean Kirmani, Brian Ichter, Dorsa Sadigh, Leonidas J. Guibas, and Fei Xia. 2024d. SpatialVLM: Endowing Vision-Language Models with Spatial Reasoning Capabilities. In CVPR. 14455–14465.
- Chen et al. (2023) Bei Chen, Fengji Zhang, Anh Nguyen, Daoguang Zan, Zeqi Lin, Jian-Guang Lou, and Weizhu Chen. 2023. CodeT: Code Generation with Generated Tests. In ICLR.
@@
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Pondé de Oliveira Pinto, Jared Kaplan, Harrison Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, Alex Ray, Raul Puri, Gretchen Krueger, Michael Petrov, Heidy Khlaaf, Girish Sastry, Pamela Mishkin, Brooke Chan, Scott Gray, Nick Ryder, Mikhail Pavlov, Alethea Power, Lukasz Kaiser, Mohammad Bavarian, Clemens Winter, Philippe Tillet, Felipe Petroski Such, Dave Cummings, Matthias Plappert, Fotios Chantzis, Elizabeth Barnes, Ariel Herbert-Voss, William Hebgen Guss, Alex Nichol, Alex Paino, Nikolas Tezak, Jie Tang, Igor Babuschkin, Suchir Balaji, Shantanu Jain, William Saunders, Christopher Hesse, Andrew N. Carr, Jan Leike, Joshua Achiam, Vedant Misra, Evan Morikawa, Alec Radford, Matthew Knight, Miles Brundage, Mira Murati, Katie Mayer, Peter Welinder, Bob McGrew, Dario Amodei, Sam McCandlish, Ilya Sutskever, and Wojciech Zaremba. 2021. Evaluating Large Language Models Trained on Code. arXiv:2107.03374 [cs.LG]
- Chen et al. (2025) Tianyu Chen, Shuai Lu, Shan Lu, Yeyun Gong, Chenyuan Yang, Xuheng Li, Md Rakib Hossain Misu, Hao Yu, Nan Duan, Peng Cheng, Fan Yang, Shuvendu K. Lahiri, Tao Xie, and Lidong Zhou. 2025. Automated Proof Generation for Rust Code via Self-Evolution. In ICLR.
- Chen et al. (2024b) Xinyun Chen, Maxwell Lin, Nathanael Schärli, and Denny Zhou. 2024b. Teaching Large Language Models to Self-Debug. In ICLR.
- Chen et al. (2022) Yueting Chen, Nick Koudas, Xiaohui Yu, and Ziqiang Yu. 2022. Spatial and Temporal Constrained Ranked Retrieval over Videos. PVLDB 15, 11 (2022), 3226–3239.
- Chen et al. (2024a) Zui Chen, Lei Cao, Sam Madden, Tim Kraska, Zeyuan Shang, Ju Fan, Nan Tang, Zihui Gu, Chunwei Liu, and Michael Cafarella. 2024a. SEED: Domain-Specific Data Curation With Large Language Models. arXiv:2310.00749 [cs.DB]
- Chen et al. (2024c) Zhenfang Chen, Rui Sun, Wenjun Liu, Yining Hong, and Chuang Gan. 2024c. GENOME: Generative Neuro-Symbolic Visual Reasoning by Growing and Reusing Modules. In ICLR.
- Cho et al. (2024) Jaemin Cho, Yushi Hu, Jason M. Baldridge, Roopal Garg, Peter Anderson, Ranjay Krishna, Mohit Bansal, Jordi Pont-Tuset, and Su Wang. 2024. Davidsonian Scene Graph: Improving Reliability in Fine-grained Evaluation for Text-to-Image Generation. In ICLR.
@@
- Dellinger et al. (2021) Justin Dellinger, Carolyn Shores, Apryle Craig, Shannon Kachel, Michael Heithaus, William Ripple, and Aaron Wirsing. 2021. Predators reduce niche overlap between sympatric prey. Oikos (12 2021).
- Deng et al. (2022) Naihao Deng, Yulong Chen, and Yue Zhang. 2022. Recent Advances in Text-to-SQL: A Survey of What We Have and What We Expect. In COLING. 2166–2187.
@@
- Feng et al. (2021) Qi Feng, Vitaly Ablavsky, and Stan Sclaroff. 2021. CityFlow-NL: Tracking and Retrieval of Vehicles at City Scale by Natural Language Descriptions. arXiv:2101.04741 [cs.CV]
@@
- Fu et al. (2019) Daniel Y. Fu, Will Crichton, James Hong, Xinwei Yao, Haotian Zhang, Anh Truong, Avanika Narayan, Maneesh Agrawala, Christopher Ré, and Kayvon Fatahalian. 2019. Rekall: Specifying Video Events using Compositions of Spatiotemporal Labels. arXiv:1910.02993 [cs.DB]
- Gao et al. (2024) Dawei Gao, Haibin Wang, Yaliang Li, Xiuyu Sun, Yichen Qian, Bolin Ding, and Jingren Zhou. 2024. Text-to-SQL Empowered by Large Language Models: A Benchmark Evaluation. PVLDB 17, 5 (2024), 1132–1145.
@@
- Hammer et al. (2020) Patrick Hammer, Tony Lofthouse, Enzo Fenoglio, Hugo Latapie, and Pei Wang. 2020. A Reasoning Based Model for Anomaly Detection in the Smart City Domain. In IntelliSys (AISC, Vol. 1251). 144–159.
@@
- He et al. (2024) Dong He, Jieyu Zhang, Maureen Daum, Alexander Ratner, and Magdalena Balazinska. 2024. MaskSearch: Querying Image Masks at Scale. arXiv:2305.02375 [cs.DB]
@@
- Hinton et al. (2015) Geoffrey Hinton, Oriol Vinyals, and Jeff Dean. 2015. Distilling the Knowledge in a Neural Network. arXiv:1503.02531 [stat.ML]
@@
- Hsieh et al. (2023a) Cheng-Yu Hsieh, Sibei Chen, Chun-Liang Li, Yasuhisa Fujii, Alexander J. Ratner, Chen-Yu Lee, Ranjay Krishna, and Tomas Pfister. 2023a. Tool Documentation Enables Zero-Shot Tool-Usage with Large Language Models. arXiv:2308.00675 [cs.CL]
@@
- Kakkar et al. (2024) Gaurav Tarlok Kakkar, Jiashen Cao, Aubhro Sengupta, Joy Arulraj, and Hyesoon Kim. 2024. Hydro: Adaptive Query Processing of ML Queries. arXiv:2403.14902 [cs.DB]
@@
- Karimi et al. (2021) Mohammad Reza Karimi, Nezihe Merve Gürel, Bojan Karlas, Johannes Rausch, Ce Zhang, and Andreas Krause. 2021. Online Active Model Selection for Pre-trained Classifiers. In AISTATS, Vol. 130. 307–315.
- Katsogiannis-Meimarakis and Koutrika (2023) George Katsogiannis-Meimarakis and Georgia Koutrika. 2023. A survey on deep learning approaches for text-to-SQL. VLDB J. 32, 4 (2023), 905–936.
@@
- Kittivorawong et al. (2024) Chanwut Kittivorawong, Yongming Ge, Yousef Helal, and Alvin Cheung. 2024. Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations. PVLDB 17, 9 (2024), 2136–2148.
@@
- Krishna et al. (2019) Ranjay Krishna, Vincent S. Chen, Paroma Varma, Michael S. Bernstein, Christopher Ré, and Li Fei-Fei. 2019. Scene Graph Prediction With Limited Labels. In ICCV. 2580–2590.
@@
- Lahiri et al. (2023) Shuvendu K. Lahiri, Sarah Fakhoury, Aaditya Naik, Georgios Sakkas, Saikat Chakraborty, Madanlal Musuvathi, Piali Choudhury, Curtis von Veh, Jeevana Priya Inala, Chenglong Wang, and Jianfeng Gao. 2023. Interactive Code Generation via Test-Driven User-Intent Formalization. arXiv:2208.05950 [cs.SE]
@@
- Ma et al. (2024) Zixian Ma, Weikai Huang, Jieyu Zhang, Tanmay Gupta, and Ranjay Krishna. 2024. m &m’s: A Benchmark to Evaluate Tool-Use for multi-step multi-modal Tasks. In ECCV, Vol. 15068. 18–34.
@@
- Mitra et al. (2024) Chancharik Mitra, Brandon Huang, Trevor Darrell, and Roei Herzig. 2024. Compositional Chain-of-Thought Prompting for Large Multimodal Models. In CVPR. 14420–14431.
@@
- Qiao et al. (2024) Wei Qiao, Tushar Dogra, Otilia Stretcu, Yu-Han Lyu, Tiantian Fang, Dongjin Kwon, Chun-Ta Lu, Enming Luo, Yuan Wang, Chih-Chun Chia, Ariel Fuxman, Fangzhou Wang, Ranjay Krishna, and Mehmet Tek. 2024. Scaling Up LLM Reviews for Google Ads Content Moderation. In WSDM. 1174–1175.
@@
- Romero et al. (2023) Francisco Romero, Caleb Winston, Johann Hauswald, Matei Zaharia, and Christos Kozyrakis. 2023. Zelda: Video Analytics using Vision-Language Models. arXiv:2305.03785 [cs.DB]
@@
- Rozière et al. (2024) Baptiste Rozière, Jonas Gehring, Fabian Gloeckle, Sten Sootla, Itai Gat, Xiaoqing Ellen Tan, Yossi Adi, Jingyu Liu, Tal Remez, Jérémy Rapin, Artyom Kozhevnikov, Ivan Evtimov, Joanna Bitton, Manish Bhatt, Cristian Canton-Ferrer, Aaron Grattafiori, Wenhan Xiong, Alexandre Défossez, Jade Copet, Faisal Azhar, Hugo Touvron, Louis Martin, Nicolas Usunier, Thomas Scialom, and Gabriel Synnaeve. 2024. Code Llama: Open Foundation Models for Code. arXiv:2308.12950 [cs.CL]
@@
- Schlining et al. (2013) Kyra Schlining, Susan von Thun, Linda Kuhnz, Brian Schlining, Lonny Lundsten, Nancy Jacobsen Stout, Lori Chaney, and Judith Connor. 2013. Debris in the deep: Using a 22-year video annotation database to survey marine litter in Monterey Canyon, central California, USA. Deep Sea Research Part I: Oceanographic Research Papers 79 (2013), 96–105.
@@
- Toubal et al. (2024) Imad Eddine Toubal, Aditya Avinash, Neil Gordon Alldrin, Jan Dlabal, Wenlei Zhou, Enming Luo, Otilia Stretcu, Hao Xiong, Chun-Ta Lu, Howard Zhou, Ranjay Krishna, Ariel Fuxman, and Tom Duerig. 2024. Modeling Collaborator: Enabling Subjective Vision Classification with Minimal Human Effort via LLM Tool-Use. In CVPR. 17553–17563.
- Wan et al. (2024) Zhongwei Wan, Xin Wang, Che Liu, Samiul Alam, Yu Zheng, Jiachen Liu, Zhongnan Qu, Shen Yan, Yi Zhu, Quanlu Zhang, Mosharaf Chowdhury, and Mi Zhang. 2024. Efficient Large Language Models: A Survey. Trans. Mach. Learn. Res. 2024 (2024).
@@
- Wang et al. (2025a) Huichen Will Wang, Larry Birnbaum, and Vidya Setlur. 2025a. Jupybara: Operationalizing a Design Space for Actionable Data Analysis and Storytelling with LLMs. arXiv:2501.16661 [cs.HC]
- Wang et al. (2025b) Huichen Will Wang, Mitchell Gordon, Leilani Battle, and Jeffrey Heer. 2025b. DracoGPT: Extracting Visualization Design Preferences from Large Language Models. IEEE Trans. Vis. Comput. Graph. 31, 1 (2025), 710–720.
- Wang et al. (2024) Jiayu Wang, Yifei Ming, Zhenmei Shi, Vibhav Vineet, Xin Wang, Sharon Li, and Neel Joshi. 2024. Is A Picture Worth A Thousand Words? Delving Into Spatial Reasoning for Vision Language Models. In NeurIPS.
@@
- Wu et al. (2023) Qingyun Wu, Gagan Bansal, Jieyu Zhang, Yiran Wu, Beibin Li, Erkang Zhu, Li Jiang, Xiaoyun Zhang, Shaokun Zhang, Jiale Liu, Ahmed Hassan Awadallah, Ryen W White, Doug Burger, and Chi Wang. 2023. AutoGen: Enabling Next-Gen LLM Applications via Multi-Agent Conversation. arXiv:2308.08155 [cs.AI]